OpenStudy (anonymous):

x - 2y = 1 x + 3y + 1z = 0 2x - 2z = 18 is the answer (1, 1, 1)?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

What brought you to (1,1,1)??

OpenStudy (anonymous):

well first you have to find the 2 new equations and i got 5y+1z=-1 and -6y-4z=18

OpenStudy (anonymous):

The second one is 1-2y-z=9 therefore -2y-z=8

OpenStudy (anonymous):

where did you get 1-2y-z=9?